Hi, I have done this problem in solving equation:
2^2 – x – 1 = 0
This is how I did the problem
Factor the equation: (2x + 1) (x – 1) = 0
Factor separately to equal 0: 2x + 1= 0 x – 1 = 0
2x = -1 x = 1
x = - ½
My question is how to check the answer, and to go back to original equation?
I hope some one can help me please I am not sure about how to go back to the original equation. Thank you very much.

Answer by stanbon(75887)

2x^2 – x – 1 = 0
2x^2-2x+x-1 = 0
2x(x-1)+(x-1) = 0
(x-1)(2x+1) = 0
x-1 = 0 or 2x+1 = 0
x = 1 or x = -1/2
----------------------
Checking:
x=1:
2*1^2 - 1 -1 = 0
2-1-1 = 0
0=0
----------
x=-1/2
2(-1/2)^2-(-1/2)-1 = 0
2(1/4) + 1/2 - 1 = 0
1/2 + 1/2 - 1 = 0
1 - 1 = 0
0 = 0
===============
Cheers,
Stan H.